OpenStudy (anonymous):

Mike surveyed the students at his school to determine which activity they would like to take part in. The bowling alley offered a 20 percent discount if 300 students of the school's 560 would sign up for the trip. Mike surveyed a random sample and asked which activity they would prefer. Bowling-60 Skating-90 Attend a play-15 Swimming-35 Based on the data in the survey,how many students are likely to choose bowling? A. 98 B. 42 C. 168 D. 252 Will they get the discount offered by the bowling alley?

OpenStudy (anonymous):

@Firejay5

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

@ash2326

OpenStudy (firejay5):

I think it's D. 252, but I am not for sure

OpenStudy (anonymous):

How did you get that; do you mind explaining?

OpenStudy (firejay5):

I am not that good at Random Samples

OpenStudy (anonymous):

Oh, ok. I'll just put 252.

OpenStudy (firejay5):

Do you trust me

OpenStudy (anonymous):

Will they get the discount offered by the bowling alley?

OpenStudy (anonymous):

And yes, sure, I trust you.

OpenStudy (firejay5):

They won't

OpenStudy (anonymous):

Okay.

OpenStudy (firejay5):

because only 60 students want to go

OpenStudy (anonymous):

Alright. Thanks!

OpenStudy (firejay5):

well wait a second

OpenStudy (anonymous):

Okay.

OpenStudy (firejay5):

Add each of the answers A, B, C, and D to 6 and see if it equals 300 or close to it

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

560.

OpenStudy (firejay5):

252 + 60 = 312 students, so they will have the discount

OpenStudy (anonymous):

Oh ok/

OpenStudy (firejay5):

Why did you say 560

OpenStudy (anonymous):

so they will

OpenStudy (anonymous):

I added 98+42+168+252

OpenStudy (anonymous):

NVM though.

OpenStudy (firejay5):

300 students of the 560 students will get the 20% discount

OpenStudy (anonymous):

Ah. So no?

OpenStudy (firejay5):

252 + 60 = 312, so yes

OpenStudy (anonymous):

Oh. Okay. Thanks!

OpenStudy (firejay5):

I believe so Idk

OpenStudy (tkhunny):

Wow... Sample Size is 200. This is an amazing 200/560 = 36% of the population. This gives the following naive proportions that might choose each option. Bowling-60/200 = 30% ==> 168 / 560 Skating-90/200 = 45% ==> 252 / 560 Attend a play-15/200 = 7.5% ==> 42 / 560 Swimming-35/200 = 17.5% ==> 98 / 560 Checking 168 + 252 + 42 + 98 = 560 -- Good. Note: This sort of survey and extrapolation is loaded with error potential. I called it "naive" in order to be VERY polite.

OpenStudy (firejay5):

@tkhunny It should be 300

OpenStudy (tkhunny):

Please define "it". The sample size is 200. The student body is 560. Use your naive predictive measure to see if your student body will meet the bowling alley's requirement of 300. Then asnwer the questions: 1) Based on the data in the survey, how many students are likely to choose bowling? 2) Will they get the discount offered by the bowling alley?

OpenStudy (firejay5):

560 is total students and sample size is 300 = 20% discount of students who want to go to the bowling alley

OpenStudy (tkhunny):

The sample size is NOT 300. Bowling-60 Skating-90 Attend a play-15 Swimming-35 60 + 90 + 15 + 35 = 200 The sample size is 200.

OpenStudy (firejay5):

Okay, what's the answer

OpenStudy (tkhunny):

It is already given in the discussion above. Please demonstrate your ability to understand the question by selecting the correct responses.

OpenStudy (firejay5):

B. 42 then

OpenStudy (anonymous):

I had a bad connection;sorry. But it's okay, Jay and Tk. You may now leave my post if you wish to do so:)

OpenStudy (firejay5):

why

OpenStudy (tkhunny):

If you want to attend the play, then 42 is correct. Try the bowling alley, instead.

OpenStudy (firejay5):

C. then

OpenStudy (tkhunny):

Now answer the final question. Is 168 > 300?